OpenStudy (anonymous):

Please help:Eliminate the parameter. x = t^2 + 1, y = t^2 - 1

OpenStudy (anonymous):

x = t^2 + 1, y = t^2 - 1 x-y=2

OpenStudy (anonymous):

Thank you! but that's it, is it? y = x + 2, x ≥ 1 y = x - 2, x ≥ 1 y = x^2 - 2, x ≥ 1 y = x^2 + 2, x ≥ 1

OpenStudy (anonymous):

@kittyxie1

OpenStudy (anonymous):

@EmilyFaye

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which of the options do i select? those are the only 4 it gave me

OpenStudy (anonymous):

i would say it is \(y=x-2\) for \(x\geq 1\)

OpenStudy (anonymous):

Thanks guys!